OpenStudy (anonymous):

-2(2s-6)-12=-3(9s+7)-3 solve the equation

OpenStudy (anonymous):

-2(2s-6)-12 = -3(9s+7)-3 -4s + 12 - 12 = -27s - 21 - 3 -4s = -27s - 24 23s = -24 s = 23/-24

OpenStudy (anonymous):

-1/6(x-12)-1/2(x+2)=x+3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

s=−24/23

OpenStudy (anonymous):

-1/6(x-12)-1/2(x+2)=x+3 -1/6 * x + 2 - .5x + 1 = x + 3 -2/3 * x + 3 = x + 3 -2/3 * x = x This isn't possible... unless x = 0

OpenStudy (texaschic101):

-2(2s - 6) - 12 = -3(9s + 7) - 3 -4s + 12 - 12 = -27s - 21 - 3 -4s = -27s - 24 -4s + 27s = -24 23s = -24 s = -24/23 or -1 1/23

OpenStudy (anonymous):

Translate to an equation, then solve. Nineteen less than six times a number is equal to 11 more than four times the number.